OpenStudy (anonymous):

You have 4 quarters, 3 dimes, 2 nickels, and 5 pennies. In how many ways can they be arranged in a row? (Coins of the same value cannot be distinguished from each other.)

OpenStudy (anonymous):

I am assuming I just multiply 4*3*2*5?

OpenStudy (anonymous):

How are you getting that?

OpenStudy (anonymous):

Let \(n\) be the total number of coins. Let \(k_i\) be the number of coins in the \(i\)th denomination.

OpenStudy (anonymous):

well because im trying to find out how many ways they can be arranged so i figured if I multiply them i'll get the amount of ways it can be arranged in a row

OpenStudy (anonymous):

Just to be clear\[ n = \sum k_i \]

OpenStudy (anonymous):

The number of ways to arrange it would be \[ \frac{n!}{\prod k_i!} \]

OpenStudy (anonymous):

This is a generalized version of the choose operation. The choose operation is basically \(k_1=k\) and \(k_2=n-k\). \[ \frac{n!}{k_1!k_2!} = \frac{n!}{k!(n-k)!} \]

OpenStudy (anonymous):

In our case, we have: \[ \frac{(4+3+2+5)!}{4!3!2!5!} \]

OpenStudy (anonymous):

Which is a very big number it seems... http://www.wolframalpha.com/input/?i=(4%2B3%2B2%2B5)!%2F(4!3!2!5!)

OpenStudy (anonymous):

Pretty interesting pattern when you think about it: \[\Large \frac{\left(\sum k_i\right)!}{\prod \left(k_i!\right)} \]